# Scalable and Robust Multiband Modeling of AGN Light Curves in Rubin-LSST

### v1

@@ 4.3 Beyond the Dirac Delta Transfer Function
EzTaoX models interband lags in multiband light curves using a Dirac delta transfer function, a simplified approach compared to methods that adopt more flexible transfer functions with finite width and skewed shapes (e.g., Zu et al., 2011; Starkey et al., 2016; Li et al., 2016; Donnan et al., 2021). While more sophisticated transfer functions can provide additional insights into the geometry and physical properties of AGN accretion disks (e.g., Cackett et al., 2007; Starkey et al., 2016; Fagin et al., 2024), we argue that the cadence of LSST WFD light curves is likely insufficient to resolve the detailed shape of the transfer functions. Thus, the simplified Dirac delta transfer function adopted by EzTaoX is likely adequate for lag estimation from LSST WFD light curves. Furthermore, previous investigations and the results of Section 3.2 show that even with high-cadence continuum reverberation mapping data, the use of more flexible transfer functions does not necessarily result in significantly different mean lag estimates compared to simpler methods (e.g., Fausnaugh et al., 2018; Donnan et al., 2023). An exception perhaps arises when the intrinsic transfer function is notably skewed, in which case simplified methods may underestimate the mean lag (Yu et al., 2020a; Chan et al., 2020).
@@ 4.4 LSST Variability Science Enabled by EzTaoX
EzTaoX enables reliable and simultaneous characterization of variability and interband lags for the tens of millions of AGNs to be monitored by LSST. Existing non-ML methods/tools capable of jointly characterizing stochastic variability and measuring interband lags are not scalable to LSST data volumes (e.g., Zu et al., 2011; Starkey et al., 2016; Donnan et al., 2021). While ML approaches are scalable, their reliability critically depends on the assumption that the training data accurately capture the true nature of AGN variability—an assumption that is rarely satisfied in practice. In comparison to traditional non-ML tools, EzTaoX achieves a speed increase of $100\times$ for LSST WFD light curves, and $10^{4}\times$ for LSST DDF light curves, without compromising —and in some cases improving—the accuracy of the results. Furthermore, EzTaoX offers the flexibility to incorporate more expressive GP kernels, allowing for improved modeling of intrinsic continuum variability. This combination of scalability and flexibility enables robust characterization of variability and interband lags for an unprecedented number of AGNs. A large, statistically significant sample of AGNs with well-characterized continuum variability and interband lags will facilitate detailed investigations into the fundamental drivers of AGN variability and the physical origins of continuum lags.
EzTaoX enables robust variability selection of actively accreting MBHs. Accreting MBHs in dwarf galaxies are difficult to identify due to their low luminosity relative to their host galaxy starlight (Greene et al., 2020). Nevertheless, their characteristic stochastic variability provides a powerful selection tool in the time domain (e.g., Baldassare et al., 2018; Ward et al., 2022; Burke et al., 2022; Bernal et al., 2025). By modeling dwarf galaxy image-differencing light curves as a DRW process, one can identify MBH candidates, which are expected to exhibit characteristic $\tau_{\rm DRW}$ values ranging from hours to days (Burke et al., 2021; Wang et al., 2023). Given the cadence of LSST WFD light curves, the multiband modeling capability of EzTaoX enables precise measurement of $\tau_{\rm DRW}$ on timescales from one hour to one day (see Figure 3), outperforming single-band approaches. When nightly stacked images are required to detect variability from fainter sources, the intra-night cadence in single-band LSST light curves is lost, leading to biased $\tau_{\rm DRW}$ estimates (see Figure 4). In contrast, through multiband fitting, EzTaoX recovers $\tau_{\rm DRW}$ without bias down to hourly timescales for AGNs detected in the WFD survey.
@@ 4.5 Caveats & Future Work
Variable emission from extended re-processors is not explicitly modeled in the current implementation of EzTaoX. Recent investigations suggest that reprocessed emission arising at distances much larger than the scale of the accretion disk—for example, in the BLR—can contribute to the measured continuum lags, potentially leading to overestimated accretion disk sizes relative to theoretical predictions (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018; Lira et al., 2024; Jaiswal et al., 2024; Leighly et al., 2024). The current multiband light curve model in EzTaoX assumes a single lag between any two photometric bands, and therefore lacks the capacity to disentangle multi-component lag distributions, such as reprocessed emission arising from the accretion disk versus that from much larger scales. Developing methodologies that can decompose observed light curves into distinct lagged components would be highly valuable—not only for probing the detailed geometry of AGN accretion flows, but also for advancing our understanding of the physical drivers behind AGN stochastic variability.
The GP kernels currently supported by EzTaoX are all stationary; however, AGN stochastic variability may exhibit non-stationary behavior. A notable example is the class of changing-look AGNs (see Ricci & Trakhtenbrot 2023 and references therein), which exhibit dramatic variability on the timescale of years to decades, likely driven by extreme changes in the accretion rate of the disk (e.g., Ruan et al., 2016; Noda & Done, 2018; Graham et al., 2020; Panda & Śniegowska, 2024). Such systems underscore the limitations of assuming stationarity in modeling AGN variability. Future development of EzTaoX will explore non-stationary GP kernels, in which the kernel parameters evolve with time rather than remaining fixed (e.g., Noack & Sethian, 2021).
@@ 5 Summary & Conclusion
We present EzTaoX —a new software package designed to leverage multiband AGN light curves for joint inference of intrinsic stochastic variability properties and interband continuum lags. We evaluate the performance of EzTaoX using both simulated LSST AGN light curves sampled at the cadence of the WFD survey, and multiwavelength light curves of real AGNs. The EzTaoX software is available on GitHub [^7] EzTaoX codebase: https://github.com/ywx649999311/EzTaoX. under a MIT License and version 0.1 is archived in Zenodo (Yu & Burke, 2025). The key advancements and capabilities of EzTaoX are summarized as follows:
@@
In addition to characterizing AGN stochastic variability and measuring continuum lags, EzTaoX is applicable to a variety of other time-domain astrophysical studies. For example, EzTaoX can be applied to measure BLR emission line lags from spectroscopic RM light curves (Blandford & McKee, 1982), or to detect negative lags that may arise from inward propagation of accretion rate fluctuations (Secunda et al., 2023)—both of which typically occur on timescales much longer than the continuum lags. Moreover, EzTaoX can be used to measure time delays between the multiple images of gravitationally lensed quasars, providing an important avenue for constraining the Hubble constant (Blandford & Narayan, 1992). Finally, interband lags have been shown to correlate with AGN continuum luminosity (e.g., Sergeev et al., 2005; Netzer, 2021; Guo et al., 2022; Panda et al., 2024); once properly calibrated, this lag-luminosity relation may enable the estimation of AGN physical properties using photometric time-domain data alone.
@@
We thank Maurizio Paolillo for providing comments to an early draft of this manuscript. W.Y. acknowledges support from the Dunlap Institute for Astronomy & Astrophysics at the University of Toronto. J.J.R. acknowledges support from the Canada Research Chairs (CRC) program, the NSERC Discovery Grant program, the Canada Foundation for Innovation (CFI), and the Québec Ministère de l’Économie et de l’Innovation. C.J.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303803. This material is based upon work supported by the National Science Foundation under Award No. 2303803. This research award to NSF is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. R.J.A., F.E.B, and S.S.S. gratefully acknowledge support from ANID: CATA BASAL FB210003 (R.J.A., F.E.B., S.S.S.); Millennium Science Initiative AIM23-0001 (F.E.B.); and FONDECYT Regular 1231718 (R.J.A.) and 1241005 (F.E.B.). D.D. acknowledges PON R&I 2021, CUP E65F21002880003, and Fondi di Ricerca di Ateneo (FRA), linea C, progetto TORNADO. L.H.G. acknowledges financial support from ANID programs: FONDECYT Iniciación 11241477, Millennium Science Initiative Programs NCN2023-002, and AIM23-0001. D.I. and A.B.K. acknowledge funding provided by the University of Belgrade - Faculty of Mathematics (the contract 451-03-136/2025-03/200104) through the grant by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia. S.P. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.

### v2

@@ 4.3 Beyond the Dirac Delta Transfer Function
EzTaoX models interband lags in multiband light curves using a Dirac delta transfer function, a simplified approach compared to methods that adopt more flexible transfer functions with finite width and skewed shapes (e.g., zu2011; starkey2016; li2016; donnan2021). While more sophisticated transfer functions can provide additional insights into the geometry and physical properties of AGN accretion disks (e.g., Cackett2007; starkey2016; fagin2024), we argue that the cadence of LSST WFD light curves is likely insufficient to resolve the detailed shape of the transfer functions. Thus, the simplified Dirac delta transfer function adopted by EzTaoX is likely adequate for lag estimation from LSST WFD light curves. Furthermore, previous investigations and the results of Section 3.2 show that even with high-cadence continuum reverberation mapping data, the use of more flexible transfer functions does not necessarily result in significantly different mean lag estimates compared to simpler methods (e.g., Fausnaugh2018; donnan2023). An exception perhaps arises when the intrinsic transfer function is notably skewed, in which case simplified methods may underestimate the mean lag (yu2020a; Chan2020).
@@ 4.4 LSST Variability Science Enabled by EzTaoX
EzTaoX enables reliable and simultaneous characterization of variability and interband lags for the tens of millions of AGNs to be monitored by LSST. Existing non-ML methods/tools capable of jointly characterizing stochastic variability and measuring interband lags are not scalable to LSST data volumes (e.g., zu2011; starkey2016; donnan2021). While ML approaches are scalable, their reliability critically depends on the assumption that the training data accurately capture the true nature of AGN variability—an assumption that is rarely satisfied in practice. In comparison to traditional non-ML tools, EzTaoX achieves a speed increase of $100\times$ for LSST WFD light curves, and $10^{4}\times$ for LSST DDF light curves, without compromising —and in some cases improving—the accuracy of the results. Furthermore, EzTaoX offers the flexibility to incorporate more expressive GP kernels, allowing for improved modeling of intrinsic continuum variability. This combination of scalability and flexibility enables robust characterization of variability and interband lags for an unprecedented number of AGNs. A large, statistically significant sample of AGNs with well-characterized continuum variability and interband lags will facilitate detailed investigations into the fundamental drivers of AGN variability and the physical origins of continuum lags.
EzTaoX enables robust variability selection of actively accreting MBHs. Accreting MBHs in dwarf galaxies are difficult to identify due to their low luminosity relative to their host galaxy starlight (greene2020). Nevertheless, their characteristic stochastic variability provides a powerful selection tool in the time domain (e.g., baldassare2018; ward2022; burke2022; bernal2025). By modeling dwarf galaxy image-differencing light curves as a DRW process, one can identify MBH candidates, which are expected to exhibit characteristic $\tau_{\rm DRW}$ values ranging from hours to days (burke2021; Wang2023a). Given the cadence of LSST WFD light curves, the multiband modeling capability of EzTaoX enables precise measurement of $\tau_{\rm DRW}$ on timescales from one hour to one day (see Figure 3), outperforming single-band approaches. When nightly stacked images are required to detect variability from fainter sources, the intra-night cadence in single-band LSST light curves is lost, leading to biased $\tau_{\rm DRW}$ estimates (see Figure 4). In contrast, through multiband fitting, EzTaoX recovers $\tau_{\rm DRW}$ without bias down to hourly timescales for AGNs detected in the WFD survey.
@@ 4.5 Caveats & Future Work
Variable emission from extended re-processors is not explicitly modeled in the current implementation of EzTaoX. Recent investigations suggest that reprocessed emission arising at distances much larger than the scale of the accretion disk—for example, in the BLR—can contribute to the measured continuum lags, potentially leading to overestimated accretion disk sizes relative to theoretical predictions (e.g., korista2001; mchardy2018; cackett2018; lawther2018; Lira2024; Jaiswal2024; Leighly2024). The current multiband light curve model in EzTaoX assumes a single lag between any two photometric bands, and therefore lacks the capacity to disentangle multi-component lag distributions, such as reprocessed emission arising from the accretion disk versus that from much larger scales. Developing methodologies that can decompose observed light curves into distinct lagged components would be highly valuable—not only for probing the detailed geometry of AGN accretion flows, but also for advancing our understanding of the physical drivers behind AGN stochastic variability.
The GP kernels currently supported by EzTaoX are all stationary; however, AGN stochastic variability may exhibit non-stationary behavior. A notable example is the class of changing-look AGNs (see Ricci2023 and references therein), which exhibit dramatic variability on the timescale of years to decades, likely driven by extreme changes in the accretion rate of the disk (e.g., Ruan2016; Noda2018; graham2020; Panda2024b). Such systems underscore the limitations of assuming stationarity in modeling AGN variability. Future development of EzTaoX will explore non-stationary GP kernels, in which the kernel parameters evolve with time rather than remaining fixed (e.g., Noack2021).
@@ 5 Summary & Conclusion
We present EzTaoX —a new software package designed to leverage multiband AGN light curves for joint inference of intrinsic stochastic variability properties and interband continuum lags. We evaluate the performance of EzTaoX using both simulated LSST AGN light curves sampled at the cadence of the WFD survey, and multiwavelength light curves of real AGNs. The EzTaoX software is available on GitHub [^6] EzTaoX codebase: https://github.com/ywx649999311/EzTaoX. under a MIT License and version 0.1 is archived in Zenodo (eztaox). The key advancements and capabilities of EzTaoX are summarized as follows:
@@
In addition to characterizing AGN stochastic variability and measuring continuum lags, EzTaoX is applicable to a variety of other time-domain astrophysical studies. For example, EzTaoX can be applied to measure BLR emission line lags from spectroscopic RM light curves (blandford1982), or to detect negative lags that may arise from inward propagation of accretion rate fluctuations (Secunda2023)—both of which typically occur on timescales much longer than the continuum lags. Moreover, EzTaoX can be used to measure time delays between the multiple images of gravitationally lensed quasars, providing an important avenue for constraining the Hubble constant (Blandford1992). Finally, interband lags have been shown to correlate with AGN continuum luminosity (e.g., sergeev2005; netzer2021; guo2022; Panda2024a); once properly calibrated, this lag-luminosity relation may enable the estimation of AGN physical properties using photometric time-domain data alone.
@@
We thank the anonymous referee for the thoughtful comments that helped us improve the paper. We thank Maurizio Paolillo and Thaisa Storchi Bergmann for providing comments to an early draft of this manuscript. W.Y. acknowledges support from the Dunlap Institute for Astronomy & Astrophysics at the University of Toronto. J.J.R. acknowledges support from the Canada Research Chairs (CRC) program, the NSERC Discovery Grant program, the Canada Foundation for Innovation (CFI), and the Québec Ministère de l’Économie et de l’Innovation. C.J.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303803. This material is based upon work supported by the National Science Foundation under Award No. 2303803. This research award to NSF is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. R.J.A., F.E.B, and S.S.S. gratefully acknowledge support from ANID: CATA BASAL FB210003 (R.J.A., F.E.B., S.S.S.); Millennium Science Initiative AIM23-0001 (F.E.B.); and FONDECYT Regular 1231718 (R.J.A.) and 1241005 (F.E.B.). D.D. acknowledges PON R&I 2021, CUP E65F21002880003, and Fondi di Ricerca di Ateneo (FRA), linea C, progetto TORNADO. L.H.G. acknowledges financial support from ANID programs: FONDECYT Iniciación 11241477, Millennium Science Initiative Programs NCN2023-002, and AIM23-0001. D.I. and A.B.K. acknowledge funding provided by the University of Belgrade - Faculty of Mathematics (the contract 451-03-136/2025-03/200104) through the grant by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia. S.P. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.